OpenStudy (anonymous):

-5 times the square of x + the square of 9x

OpenStudy (saifoo.khan):

\[-5 \times x + \sqrt{9x}\]

OpenStudy (saifoo.khan):

it can't be that.

OpenStudy (anonymous):

they won't have a times sign, htey just have is beside it

OpenStudy (anonymous):

Assuming that's sqrt not square. \[-5\sqrt{x}+\sqrt{9x}=-5\sqrt{x}+\sqrt{3^2*x}\] \[-5\sqrt{x}+3\sqrt{x}=-2\sqrt{x}\]

OpenStudy (anonymous):

\[-5x^2+(9x)^2\]?

OpenStudy (anonymous):

that's what I understand too Satellite - wolfgirl ?

OpenStudy (anonymous):

\[-5\sqrt{x} + \sqrt{9x}\] \[-5\sqrt{x} +3\sqrt{x}\] \[-2\sqrt{x}\]

OpenStudy (anonymous):

thts hard to explain fiddle, Mohoney and Osirisis got it right

OpenStudy (anonymous):

hard to explain ? All we were asking was for you to confirm if the question was asking about square roots or squares ? Most ppl answering here had that doubt :)

OpenStudy (anonymous):

its about simplifying equations and thinking about perfect squares when working with square roots

OpenStudy (anonymous):

yeah it's the square root I think